The speed of rain A travels t10 miles slower than train B. Train A traveled 200 miles in the same time it took train B to go 250 miles. find the speed for train A and B. Is the formula for this d=k-r2? Or how would I figure this out.
Answer by mananth(16949)

let speed of train B be x mph
distance traveled by train B = 250 miles.
time taken by train B = 250/x
..
Speed of train A =x+10
distance traveled by train A = 200/x-10
..
The times taken are equal.
250/x = 200/x-10
250(x-10)=200x
250x-2500=200x
50x= 2500
x=50 mph --- speed of train B
Train A travels 40 mph